OpenStudy (anonymous):

The radius of a circle on a basketball court is 6 feet. what are the approximate circumference and area of the circle

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (jdoe0001):

\(\bf \textit{area of a circle}=\pi r^2\qquad circumference=2\pi r\qquad r=radius\)